Kesha has collected the autographs of 9 more television celebrities than she has of movie stars. each tv celebrity autograph is worth $75 and each movie star autograph is worth $250. her colection is valued at $2,300.
how many Tv celebrity autographs does she have?
how many of movie star autographs does she have?

Answer by chitra(359)

Let "x" represent the autograph of the movie stars and "y" represent the autograph of the tv celebrities.
So according to the given data, she has 9 more autographs of the tv celebs than the movie stars.
So y = x + 9

Now let us wirte the equation

250x + 75y = 2300

Substitute for "y" we get:

250x + 75(x + 9) = 2300

250x + 75x + 675 = 2300

325x = 1625

x = 5

Pluggin back we get the value of y.

y = x + 9

y = 5 + 9

y = 14

Thus, the solution.